# Boundless Byte Pair Encoding: Breaking the Pre-tokenization Barrier

### v1

@@ Abstract
Pre-tokenization, the initial step in many modern tokenization pipelines, segments text into smaller units called pretokens, typically splitting on whitespace and punctuation. While this process encourages having full, individual words as tokens, it introduces a fundamental limitation in most tokenization algorithms such as Byte Pair Encoding (BPE). Specifically, pre-tokenization causes the distribution of tokens in a corpus to heavily skew towards common, full-length words. This skewed distribution limits the benefits of expanding to larger vocabularies, since the additional tokens appear with progressively lower counts. To overcome this barrier, we propose BoundlessBPE, a modified BPE algorithm that relaxes the pretoken boundary constraint. Our approach selectively merges two complete pretokens into a larger unit we term a superword. Superwords are not necessarily semantically cohesive. For example, the pretokens ␣of and ␣the might be combined to form the superword ␣of␣the. This merging strategy results in a substantially more uniform distribution of tokens across a corpus than standard BPE, and compresses text more effectively, with an approximate 20% increase in bytes per token.
@@ 1 Introduction
Pre-tokenization is a crucial step in preparing text for language models, helping to align token boundaries to meaningful linguistic units. A document is first broken into chunks called pretokens using a regular expression, [^1] See Section 3.5 and Appendix A for a discussion of the particular pre-tokenizer regular expression we used. which are then tokenized separately. Each pretoken may be tokenized into two or more subword tokens, or used as a single token that exactly matches the entire pretoken. Many common words end up as a single token (Reddy et al., 2025), more than 90% according to our analysis (Section 2). Under such circumstances, differences in the tokenizer itself can only manifest in the small percentage of remaining pretokens. This high overlap in resulting tokens shared across different tokenizers can explain why tokenizers have been found to perform quite similarly on downstream tasks, with no statistically superior approach (Schmidt et al., 2024). Selecting a larger vocabulary size tends to exacerbate this issue, since the lower-frequency tokens added to the vocabulary later in the training process do not substantially change the resulting token distribution in the training corpus.
@@
With the goal of obtaining a more uniform distribution of tokens, we propose a modification to the standard pre-tokenization approach that still allows us to retain the benefits of contextual cohesion. Specifically, we introduce superwords, tokens composed of an $n$-gram of words, to supplement subwords and words in the vocabulary. The notion of superwords extends beyond semantically cohesive units like New York City often identified through metrics like Pointwise Mutual Information (PMI; Fano & Wintringham, 1961; Church & Hanks, 1989). Instead, they serve to distribute the most common words like ␣the into a range of common $n$-grams like ␣the␣car and ␣the␣house, thereby lowering the frequency of the most common tokens in a corpus. As we will demonstrate, a substantial number of such common $n$-grams is found by our approach, enabling more effective utilization of larger vocabularies.
In Section 3, we introduce BoundlessBPE, an extension to the widely adopted Byte Pair Encoding (BPE) algorithm (Sennrich et al., 2016; Gage, 1994). The key modification involves incorporating the concept of a supermerge. During the tokenizer training process, we permit the merging of two adjacent pretokens if each is represented by a single token, in which case, the resulting merge is considered a superword. ␣the␣car, when the co-occurrence counts are sufficiently high. The supermerge operation lends itself to seamless integration into the standard BPE training procedure: at each step of training, the merge or permissible supermerge that occurs the most frequently is performed. We also implement a variation of the deletion technique from PickyBPE (Chizhov et al., 2024) to remove low-frequency intermediate tokens from the vocabulary. These deletions contribute to a more uniform distribution at the lower end of the frequency spectrum by freeing up space taken by intermediate tokens for use by more common tokens.
The efficacy of our proposed approach is presented in Section 4, where we use an evaluation corpus to compare the token frequency distributions produced by BoundlessBPE and several commonly-used baseline tokenizers. The results demonstrate that BoundlessBPE yields a more uniform token distribution, consequently achieving a minimum of a 21% increase in Rényi efficiency (Zouhar et al., 2023) compared to the baselines. Since our method exhibits improved compression, with at least a 19.7% increase in overall bytes per token, it requires fewer tokens for language model inference.
@@ 2 Limitations of pre-tokenizers
Due to the Zipfian nature of text, where a few words occur with very high frequency and many words occur rarely, the pretokens of common words are often efficiently represented as a single token that exactly matches the entire pretoken. We used a 5GB portion of MiniPILE (Kaddour, 2023) as our out-of-sample evaluation corpus to illustrate this. [^2] We train tokenizers on the first 170,721 documents of MiniPile, totalling 1GB. The remaining 829,279 documents form the 5GB out-of-sample evaluation corpus used throughout this paper. Using the widely adopted BPE, WordPiece  (Wu et al., 2016; Schuster & Nakajima, 2012) and UnigramLM (Kudo, 2018) tokenizers, [^3] We use the Hugging Face implementations: https://github.com/huggingface/tokenizers. with vocabulary sizes of 40,960, 65,536, 98,304, and 131,072, we calculated the proportion of pretokens that were ultimately represented by a single token. As shown in Figure 1, BPE and WordPiece, both of which are bottom-up merge-based tokenizers, achieve a single-token representation for well over 90% of pretokens, while UnigramLM, a top-down ablation-based tokenizer, achieves 82-87%. This highlights how these tokenization methods effectively handle the most frequent words in the vocabulary, and it is a direct consequence of the Zipfian distribution inherent in natural language.
@@ 3 BoundlessBPE
We introduce BoundlessBPE, a tokenization algorithm that allows adjacent pretokens, each represented as a single token, to be merged into a superword token. For the Shakespearean example above, our tokenization training process will continue beyond pretoken boundaries, yielding the tokens: [^4] BoundlessBPE examples throughout this paper are for a training run culminating with a vocabulary of 131,072.
@@ 3.1 Tokenizer training
can be combined into the superword [ of the’], which would consist of a single token. A superword can subsequently be merged with other fully-merged pretokens into longer superwords. [^5] To avoid combining different pretoken categories, such as digits and words, we restrict the merging process to pretokens match the byte regex rb"^[ _’a-zA-Z]*[a-zA-Z][ _’a-zA-Z]*\$". This regex matches words with optional initial spaces, contractions, and snake case variables. See Section 3.5 and Appendix A for a more detailed explanation of the pre-tokenization regex. Note that numbers are already segmented into groups of 3 in a right-to-left manner, following the suggestion of Singh & Strouse (2024), so we avoid combining these digit groups into longer superwords to preserve their intended format. Furthermore, as discussed in Section A.1, we opted not to combine punctuation with words. Supermerges often involve pairs of common words, such as of the’, in the’, to the’, on the’, and and the’. One consequence of this behavior is a reduction in the count of the most common tokens such as the’ and python’ of’, as some of their occurrences are allocated to various superwords.
@@ 3.2 Inference procedure
Once a tokenizer has been trained, it can be used to tokenize a new document, a process often referred to as segmentation or inference. A trained BoundlessBPE tokenizer consists of separate dictionaries for regular merge rules, supermerge rules, and regular deletion rules. Each rule has a unique index representing the order in which it was added during the training process. Each step of inference involves identifying possible operations that can be performed given the current tokenization of the document. Each pair of adjacent tokens in each pretoken is compared to the merge rules. Each pair of adjacent pretokens containing single tokens is compared to the supermerge rules. Finally, each individual token is compared to the deletion rules. The operation with the smallest index, meaning it was the first to be added during the training process, is performed everywhere it occurs within the document. Figure 2 gives an example of this tokenization process for the phrase ’Tip of the hat’:
@@
The data is initialized on line 1 with a list of 4 pretokens, each containing a list of individual bytes as the initial tokens. Line 2 shows the results of the first merge of ( , ’t’), which had the lowest index of any operation. This process of selecting the lowest index merge, supermerge, or deletion continues until line 7 where [ of’] and [ the’] appear next to each other, and each is a single token. At this point, supermerge becomes a valid option. On line 8, a regular merge ( , ’h’) is performed next due to a lower index, but then the first supermerge creating [ of the’] happens on line 9. The tokenization process concludes on line 12 after the application of a few more merges.
@@ 3.3 Efficient training implementation
In the representation shown in Section 3.2, a document is segmented into pretokens, with each pretoken containing one or more tokens. While conceptually straightforward and directly applicable to tokenizer training, this representation is computationally inefficient. See Appendix B for a description of several techniques used to speed up training time. We will release an open source version of the BoundlessBPE training and inference code. [^7] URL to be published after anonymous review.
@@ 3.4 Training dynamics
Figure 4 shows the logarithm of the count of each selected merge or supermerge during the training process for BoundlessBPE, along with the count of each selected merge for standard BPE and PickyBPE, each employing two distinct pre-tokenization regular expressions. The addition of supermerges as an available choice allows the counts to decrease at a slower rate compared to the baseline methods. As we will demonstrate, this results in a more uniform token distribution characterized by a tail of tokens with higher frequencies.
@@
Figure 4 shows the fraction of regular merges and supermerges across each successive group of 8,192 merges. A supermerge cannot form until both candidate pretokens are a single tokens. For example, [ of’] and [ the’] need to be represented as single tokens of’ and the’, respectively, before they are eligible for a supermerge. As a result, early in the process regular merges constitute a majority of operations, and over time the proportion of supermerges grows. Supermerges were 31.3% of total merges over the entire range to a vocabulary size of 131,072. In the rightmost interval, supermerges constitute 44.4% of operations. Thus, the BoundlessBPE tokenizer uses a substantial number of superwords, breaking through the pre-tokenization barrier.
@@ 3.5 Superwords enable improved pre-tokenization
The BOUNDLESS_PATTERN regex in Figure 5 contains a number of improvements over existing pre-tokenization regex patterns. [^8] We discuss existing regular expressions and compare them to ours in Appendix A. One notable improvement is the better handling of names in code, which is directly enabled by the supermerges in BoundlessBPE. The variable name ’XMLHttpRequest’ is composed of three sub-components: ’XML’, ’Http’, and ’Request’, which can be identified by capitalization conventions. The B-2 to B-6 branches of Figure 5 work together to break variable and function names into smaller pretokens based on capitalization. These sub-components can then be recombined via supermerges as their co-occurrence counts warrant. The example ’XMLHttpRequest snake_case camelCase CONSTANT’ is pre-tokenized as:
@@
Using the the commonly used GPT2_PATTERNfootnotemark: pre-tokenization segments on the full names, unable to utilize the prior knowledge encoded in the sub-components:
@@
Conversely, using BOUNDLESS_PATTERN with BPE would always result in the final tokens being the individual sub-components. With the supermerges BoundlessBPE provides, it is possible to both align with sub-components and to recombine some of the full names.
@@ 4 Token distribution
We have seen that pre-tokenization influences more than 90% of the tokens in the vocabulary, which makes it difficult for algorithms like BPE to alter the distribution of token occurrences. However, supermerges offer a means to overcome this barrier.
@@
The left column of Figure 6 shows the log counts for each token, sorted in descending order of frequency along the $x$-axis, in our evaluation corpus. We observe that the tail of BoundlessBPE’s token frequency chart is substantially higher than that of the three baseline tokenizers for two vocabulary sizes, [^9] See Figure 11 in Appendix C for two additional vocabulary sizes. and it has far fewer tokens that don’t appear in the evaluation corpus (plotted at zero). This is due to a combination of higher counts for final merges mentioned above and the deletion of intermediate tokens, which makes room for additional useful tokens. The right column of Figure 6 presents a zoomed-in view of this same measure for the 250 tokens with the largest count in each vocabulary, showing that supermerges have indeed succeeded in reducing the counts of overly-general tokens.
Figure 7 shows the fraction of the vocabulary that is used at least once on the same evaluation corpus. Over 97.5% of tokens in BoundlessBPE’s vocabulary are found in the evaluation corpus, across all four vocabulary sizes, while baseline tokenizers’ vocabularies include between 82-95%.
@@
Finally, we quantify the token distributions over the corpus using two metrics: (1) the Rényi efficiency metric (Zouhar et al., 2023), which indicates how uniform a token distribution is; and (2) compression rate of the evaluation corpus. Figure 9 shows that the BoundlessBPE Rényi efficiency is at least 21% above that of the baselines, using the $\alpha=2.5$ value recommended by Zouhar et al. (2023). Additional analysis of the Rényi efficiency is provided in Appendix E.
@@
Figure 9 shows that BoundlessBPE provides at least a 19.7% increase in overall bytes per token for the 5GB evaluation corpus compared to baseline tokenizers. Compression rate continues to increase at larger vocabularies, indicating it is able to effectively use the additional vocabulary space. While the effect of compression on downstream performance is unclear (Goldman et al., 2024; Ali et al., 2024; Schmidt et al., 2024; Gallé, 2019), having more bytes per token can speed up language model inference, as fewer tokens are needed to process or generate a given text.

### v2

@@ Abstract
Pre-tokenization, the initial step in many modern tokenization pipelines, segments text into smaller units called pretokens, typically splitting on whitespace and punctuation. While this process encourages having full, individual words as tokens, it introduces a fundamental limitation in most tokenization algorithms such as Byte Pair Encoding (BPE). Specifically, pre-tokenization causes the distribution of tokens in a corpus to heavily skew towards common, full-length words. This skewed distribution limits the benefits of expanding to larger vocabularies, since the additional tokens appear with progressively lower counts. To overcome this barrier, we propose BoundlessBPE, a modified BPE algorithm that relaxes the pretoken boundary constraint. Our approach selectively merges two complete pretokens into a larger unit we term a superword. Superwords are not necessarily semantically cohesive. For example, the pretokens ␣of and ␣the might be combined to form the superword ␣of␣the. This merging strategy results in a substantially more uniform distribution of tokens across a corpus than standard BPE, and compresses text more effectively, with up to a 15% increase in bytes per token.
@@ 1 Introduction
Pre-tokenization is a crucial step in preparing text for language models, helping to align token boundaries to meaningful linguistic units. A document is first broken into chunks called pretokens using a regular expression, [^1] See Section 3.5 and Appendix A for a discussion of the particular pre-tokenizer regular expressions we used. which are then tokenized separately. Each pretoken may be tokenized into two or more subword tokens, or used as a single token that exactly matches the entire pretoken. Many common words end up as a single token (Reddy et al., 2025), more than 90% for the best baselines (Section˜2). Under such circumstances, differences in the tokenizer itself can only manifest in the small percentage of remaining pretokens. This high overlap in resulting tokens shared across different tokenizers can explain why tokenizers have been found to perform quite similarly on downstream tasks, with no statistically superior approach (Schmidt et al., 2024). Selecting a larger vocabulary size tends to exacerbate this issue, since the lower-frequency tokens added to the vocabulary later in the training process do not substantially change the resulting token distribution in the training corpus.
@@
With the goal of obtaining a more uniform distribution of tokens, we propose a modification to the standard pre-tokenization approach that still allows us to retain the benefits of contextual cohesion. Specifically, we introduce superwords, tokens composed of an $n$-gram of words, to supplement subwords and words in the vocabulary. The notion of superwords extends beyond semantically cohesive units like New␣York␣City often identified through metrics like Pointwise Mutual Information (PMI; Fano & Wintringham, 1961; Church & Hanks, 1989). Instead, they serve to distribute the most common words like ␣the into a range of common $n$-grams like ␣the␣car and ␣the␣house, thereby lowering the frequency of the most common tokens in a corpus. As we will demonstrate, a substantial number of such common $n$-grams is found by our approach, enabling more effective utilization of larger vocabularies.
In Section˜3, we introduce BoundlessBPE, an extension to the widely adopted Byte Pair Encoding (BPE) algorithm (Sennrich et al., 2016; Gage, 1994). The key modification involves incorporating the concept of a supermerge. During the tokenizer training process, we permit the merging of two adjacent pretokens if each is represented by a single token, in which case, the resulting merge is considered a superword. The supermerge operation lends itself to seamless integration into the standard BPE training procedure: at each step of training, the merge or permissible supermerge that occurs the most frequently is performed. We also implement a variation of the deletion technique from PickyBPE (Chizhov et al., 2024) to remove low-frequency intermediate tokens from the vocabulary. These deletions contribute to a more uniform distribution at the lower end of the frequency spectrum by freeing up space taken by intermediate tokens for use by more common tokens.
The efficacy of our proposed approach is presented in Section˜4, where we use an evaluation corpus to compare the token frequency distributions produced by BoundlessBPE and several commonly-used baseline tokenizers. The results demonstrate that BoundlessBPE yields a more uniform token distribution, consequently achieving a 3-5% increase in Rényi efficiency (Zouhar et al., 2023) compared to the baselines. Since our method exhibits improved compression, with a 9-15% increase in overall bytes per token, it requires fewer tokens for language model inference.
@@ 2 Limitations of pre-tokenizers
Due to the Zipfian nature of text, where a few words occur with very high frequency and many words occur rarely, the pretokens of common words are often efficiently represented as a single token that exactly matches the entire pretoken. We used a 5GB portion of MiniPILE (Kaddour, 2023) as our out-of-sample evaluation corpus to illustrate this. [^2] We train tokenizers on the first 170,721 documents of MiniPile, totalling 1GB. The remaining 829,279 documents form the 5GB out-of-sample evaluation corpus used throughout this paper. Using the widely adopted BPE, WordPiece  (Wu et al., 2016; Schuster & Nakajima, 2012) and UnigramLM (Kudo, 2018) tokenizers, [^3] We use the Hugging Face implementations: https://github.com/huggingface/tokenizers. with vocabulary sizes of 40,960, 65,536, 98,304, and 131,072, we calculated the proportion of pretokens that were ultimately represented by a single token. As shown in Figure˜1, BPE and WordPiece, both of which are bottom-up merge-based tokenizers, achieve a single-token representation for 90-95% of pretokens, while UnigramLM, a top-down ablation-based tokenizer, achieves a lower range of 72-73%. This highlights how these tokenization methods effectively handle the most frequent words in the vocabulary, and it is a direct consequence of the Zipfian distribution inherent in natural language.
@@ 3 BoundlessBPE
We introduce BoundlessBPE, a tokenization algorithm that allows adjacent pretokens, each represented as a single token, to be merged into a superword token. For the Shakespearean example above, our tokenization training process will continue beyond pretoken boundaries, yielding the tokens: [^4] BoundlessBPE examples throughout this paper are with a vocabulary size of 131,072.
@@ 3.1 Tokenizer training
[…]
@@ 3.2 Inference procedure
Once a tokenizer has been trained, it can be used to tokenize a new document, a process often referred to as segmentation or inference. A trained BoundlessBPE tokenizer consists of separate dictionaries for regular merge rules, supermerge rules, and regular deletion rules. Each rule has a unique index representing the order in which it was added during the training process. Each step of inference involves identifying possible operations that can be performed given the current tokenization of the document. Each pair of adjacent tokens within each pretoken is compared to the merge rules. Each pair of adjacent pretokens containing single tokens is compared to the supermerge rules. Finally, each individual token is compared to the deletion rules. The operation with the smallest index, meaning it was the first to be added during the training process, is performed everywhere it occurs within the document. Figure˜2 gives an example of this tokenization process for the phrase ’Tip of the hat’:
@@
The data is initialized on line 1 with a list of 4 pretokens, each containing a list of individual bytes as the initial tokens. Line 2 shows the results of the first merge of ( , ’t’), which had the lowest index of any operation. This process of selecting the lowest index merge, supermerge, or deletion continues until line 7 where [ of’] and [ the’] appear next to each other, and each is a single token. At this point, a supermerge becomes a valid option. On line 8, a regular merge ( , ’h’) is performed next due to a lower index, but then the first supermerge creating [ of the’] happens on line 9. The tokenization process concludes on line 12 after the application of a few more merges.
@@ 3.3 Efficient training implementation
In the representation shown in Section˜3.2, a document is segmented into pretokens, with each pretoken containing one or more tokens. While conceptually straightforward and directly applicable to tokenizer training, this representation is computationally inefficient. See Appendix˜C for a description of several techniques used to speed up training time. We have released an open source version of the BoundlessBPE training and inference code. [^7] https://github.com/kensho-technologies/boundlessbpe
@@ 3.4 Training dynamics
Figure˜4 shows the logarithm of the count of each selected merge or supermerge during the training process for BoundlessBPE, along with the count of each selected merge for standard BPE and PickyBPE, each employing two distinct pre-tokenization regular expressions. The addition of supermerges as an available choice allows the counts to decrease at a slower rate compared to the baseline methods. As we will demonstrate, this results in a more uniform token distribution characterized by a tail of tokens with higher frequencies. The other four curves are essentially indistinguishable, which is a result of so many pretokens ending up as single tokens, as described in Section˜2.
@@
Figure˜4 shows the fraction of regular merges and supermerges across each successive group of 8,192 merges. A supermerge cannot form until both candidate pretokens are a single tokens. For example, [ of’] and [ the’] need to be represented as single tokens of’ and the’, respectively, before they are eligible for a supermerge. As a result, early in the process regular merges constitute a majority of operations, and over time the proportion of supermerges grows. Supermerges were 36.6% of total merges over the entire range to a vocabulary size of 131,072. In the rightmost interval, supermerges constitute 50.8% of operations. Thus, the BoundlessBPE tokenizer uses a substantial number of superwords, breaking through the pre-tokenization barrier.
@@ 3.5 Superwords enable improved pre-tokenization
The BOUNDLESS_PATTERN regex in Figure˜5 contains a number of improvements over existing pre-tokenization regex patterns. [^8] We discuss existing regular expressions and compare them to ours in Appendix A. One notable improvement is the better handling of names in code, which is directly enabled by the supermerges in BoundlessBPE. The variable name ’XMLHttpRequest’ is composed of three sub-components: ’XML’, ’Http’, and ’Request’, which can be identified by capitalization conventions. The B-2 to B-6 branches of Figure˜5 work together to break variable and function names into smaller pretokens based on capitalization. These sub-components can then be recombined via supermerges as their co-occurrence counts warrant. The example ’XMLHttpRequest snake_case camelCase CONSTANT’ is pre-tokenized as:
@@
Using the commonly used GPT2_PATTERNfootnotemark: pre-tokenization segments on the full names, unable to utilize the prior knowledge encoded in the sub-components:
@@
Conversely, using BOUNDLESS_PATTERN with standard BPE would always result in the final tokens being the individual sub-components. With the supermerges BoundlessBPE provides, it is possible to both align with sub-components and to recombine some of the full names.
@@ 4 Token distribution
We have seen that pre-tokenization influences more than 90% of the tokens in the vocabulary with BPE, which makes it difficult for algorithms like BPE to alter the distribution of token occurrences. However, supermerges offer a means to overcome this barrier.
@@
The left column of Figure˜6 shows the log counts for each token in our evaluation corpus, sorted in descending order of frequency along the $x$-axis. We observe that the tail of BoundlessBPE’s token frequency chart is substantially higher than that of the three baseline tokenizers [^9] BoundlessBPE (B) uses the BOUNDLESS_PATTERN regex, while BoundlessBPE (4o) uses GPT4O_PATTERN regex, as described in Appendix A. The baselines use the GPT4O_PATTERN regex. over two vocabulary sizes, [^10] See Figure 12 in Appendix B for two additional vocabulary sizes.. This is due to a combination of higher counts for final merges mentioned above and the deletion of intermediate tokens, which makes room for additional useful tokens. The right column of Figure˜6 presents a zoomed-in view of this same measure for the 250 tokens with the largest counts in each vocabulary, showing that supermerges have succeeded in reducing the counts of overly-general tokens.
Figure˜7 shows the fraction of the vocabulary that is used at least once on the same evaluation corpus. The BoundlessBPE (B) run did best across vocabulary sizes, with BoundlessBPE (4o) also doing well. [^11] Note that Figure 7 to Figure 9 use the effective vocabulary size for BoundlessBPE, subtracting the number of deletions, to make a more accurate comparison.
@@
Finally, we quantify the token distributions over the corpus using two metrics: (1) the Rényi efficiency metric (Zouhar et al., 2023) which indicates how uniform a token distribution is; and (2) compression rate of the evaluation corpus. Figure˜9 shows that BoundlessBPE (4o) achieves a Rényi efficiency at least 3% above that of the baselines. [^12] We use $\alpha=2.5$ as recommended by Zouhar et al. (2023). BoundlessBPE (B) had a lower Rényi efficiency, showing the sensitivity of this metric to the regex and value of $\alpha$. [^13] See Appendix D for results at other $\alpha$ values. There is later work pointing out limitations of Rényi efficiency in predicting downstream performance (Cognetta et al., 2024b). However, our goal was to design a system with a more uniform distribution of tokens, and it does provide a quantitative measure of the uniformity.
@@
Figure˜9 shows that BoundlessBPE (4o) provides a 9-15% increase in overall bytes per token for the 5GB evaluation corpus compared to BPE and WordPiece. The compression rate continues to increase at larger vocabularies, indicating it is able to effectively use the additional vocabulary space. While the effect of compression on downstream performance is unclear (Gallé, 2019; Goldman et al., 2024; Ali et al., 2024; Schmidt et al., 2024), having more bytes per token can speed up language model inference, as fewer tokens are needed to process or generate a given text.
5 Ablation study
BoundlessBPE combined supermerges with a new pre-tokenization regular expression, BOUNDLESS_PATTERN. We also used a different form of PickyBPE deletions, where deleted tokens were split into individual bytes before being allowed to re-merge. Our motivation was that the more aggressive decomposition could allow the deleted token to form more compressed tokens. In contrast, Chizhov et al. (2024) break a deleted token into the tokens that created it in a pairwise merge.
Vocab Size Regex Pattern PickyBPE Bytes Per Token Rényi Efficiency
40,960 Boundless none 4.268 0.4650
40,960 Boundless original 4.267 0.4650
40,960 Boundless ours 4.261 0.4651
40,960 GPT4o none 4.424 0.4925
40,960 GPT4o original 4.422 0.4925
40,960 GPT4o ours 4.416 0.4926
131,072 Boundless none 4.771 0.4048
131,072 Boundless original 4.769 0.4049
131,072 Boundless ours 4.762 0.405
131,072 GPT4o none 5.001 0.4298
131,072 GPT4o original 4.999 0.4299
131,072 GPT4o ours 4.991 0.4300
Table 1: Effect of regular expression pattern and PickyBPE style on intrinsic measures
Table˜1 gives an ablation study that disentangles these changes. [^14] See Appendix E for the same results in two additional vocabulary sizes. We consider both forms of PickyBPE deletions, as well as no deletions, and compare our BOUNDLESS_PATTERN to a modern baseline of the GPT4o regular expression. [^15] See Appendix A We see that the differences in the types of deletions had only a very small effect on either of our intrinsic measures. Still, deletions may prove important for downstream performance, since low-frequency tokens are harder to learn (Su et al., 2024; Yu et al., 2022). Substituting the regex for the GPT4o version increased bytes per token by 0.15-0.23, and increased Rényi efficiency by around 0.025. The GPT4O_PATTERN discussed in Appendix˜A contains O-1 and O-2, which allow a single non-alphanumeric character to combine with word tokens, compressing more than BOUNDLESS_PATTERN, which was designed to keep character classes more separate.